OpenStudy (anonymous):

Find the exact value of each expression without a calc.? cos(-π/6)sin(-π/4)-tanπ/4

OpenStudy (anonymous):

cos(-π/6)=-sqrt3/2 sin(-π/4)=-sqrt2/2

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

cos (-x)=cos(x) cos(pi/6)=sqrt3/2 sin(pi/4)=sqrt2/2 tan(pi/4)=1 now solve

OpenStudy (anonymous):

So i get the sol for the 3 equations separately and that's just my answer? Why are the eq stuck together then?

OpenStudy (anonymous):

multiply them

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

sqrt 6/2?

OpenStudy (anonymous):

so the answer is sqrt3/2 * (-sqrt2/2) -1

OpenStudy (anonymous):

sqrt6 - 2/2?

OpenStudy (anonymous):

that should be your answer!

OpenStudy (anonymous):

Awesome, great thanks :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good